# Graph Signal Denoising Using Regularization by Denoising and Its Parameter Estimation

### v1

@@ 3.2 PnP-ADMM
In practice, PnP-ADMM outperforms the model- and data-driven approaches for graph signal denoising yazaki2019Interpolation; cai_unrolling_2025. However, its overall optimization problem is not always explicit since $s(\cdot)$ cannot be derived from $\mathcal{D}_{\text{graph}}(\cdot)$ straightforwardly.
@@ 5 Experiments
In this section, we demonstrate the effectiveness of the proposed method through graph signal denoising for synthetic and real-world data.
@@ 5.1.1 Datasets
with a constant value $C$. In this experiment, the graph signal was generated with $N_{\text{band}}=3$ and $C=2$.
@@ 5.1.2 Comparison Methods
Methods to be compared are summarized in Table 1. They include both of model-based and data-driven methods. Hyperparameters in the model-based methods are determined from training data using Bayesian optimization optuna_2019. The network architecture of GAT is the same as rey2022Untrained. The number of iterations $K$ of the proposed method and PnP-ADMM is set to $K=10$. The internal graph denoiser $\mathcal{D}_{\text{graph}}$ for PnP-ADMM is LR. The proposed method uses two denoisers, LR or PnP-ADMM, for regularization: Their abbreviations are RED (LR) and RED (PnP), respectively. In the proposed supervised and unsupervised methods, parameters are trained using Adam DBLP:journals/corr/KingmaB14 with the learning rate $0.01$. In the proposed unsupervised method, $\sigma_{\text{N2N}}$ in (8) is randomly selected from $0,0.4\cdot\max(\mathbf{|x|})]$.
@@ 5.2 Results: Synthetic Dataset
Table 1: Overview of comparison and proposed methods.
Method type Name # of Parameters
Supervised Model-based Laplacian Regularization (LR) pang2017Graph 1
Plug-and-play ADMM (PnP) yazaki2019Interpolation 2
RED (LR) 2
RED (PnP) 3
DAU-based RED (LR-DAU) 22
RED (PnP-DAU) 33
Unsupervised Data-driven Graph Attention Network (GAT) velickovic2018Graph 8501
Untrained Graph Neural Networks (UGNN) rey2022Untrained 7550
DAU-based RED (LR-Unsupervised) 22

### v2

@@ 3.2 PnP-ADMM
In practice, PnP-ADMM outperforms the model- and data-driven approaches for graph signal denoising yazaki2019Interpolation; cai2025Unrolling. However, its overall optimization problem is not always explicit since $s(\cdot)$ cannot be derived from $\mathcal{D}_{\text{graph}}(\cdot)$ straightforwardly.
@@ 4.1.2 GAT and PnP-ADMM
Note that, a well-trained GNN for signal denoising can be readily expected to satisfy these conditions. Furthermore, it is also possible to use GNNs that are theoretically proven to satisfy them velasco2024Graph; nt2019Revisiting; jia2023Stabilizing.
We should be careful when using methods where dynamic ranges between the input and output change significantly, such as GNNs for graph classification. In this case, it is necessary to verify that the conditions are satisfied.
@@
Note that both LR and PnP internally require matrix inversion or eigendecomposition, making them computationally expensive for large-scale graphs. To address this issue, several approximation techniques could be straightforwardly employed. For instance, the matrix inversion can be efficiently approximated using methods like Chebyshev polynomial approximation hammond2011Wavelets; shuman2011Chebyshev. Furthermore, it is also possible to incorporate existing denoising methods designed for large-scale graphs zhang2023Subgraph; wu2023SGFormer, such as those based on distributed processing or graph partitioning, as the inner denoiser within our framework.
@@ 5 Experiments
In this section, we demonstrate the effectiveness of the proposed method through graph signal denoising for synthetic and real-world data [^1] Official implementation is available at https://github.com/kojima-msp/GRED_DAU..
@@ 5.1.1 Datasets
with a constant value $C$. In this experiment, the original graph signal was generated with $N_{\text{band}}=3$ and $C=2$, and scaled to $[0,100]$.
@@
Table 1: Overview of comparison and proposed methods. “S:” and “U:” denote “Supervised” and “Unsupervised,” respectively.
Method type Name # of Parameters
S: Model-based Laplacian Regularization (LR) pang2017Graph 1
Plug-and-play ADMM (PnP) yazaki2019Interpolation 2
RED (LR) 2
RED (PnP) 3
S: DAU-based RED (LR-DAU) 22
RED (PnP-DAU) 33
U: Data-driven Graph Attention Network (GAT) velickovic2018Graph 8501
Untrained Graph Neural Networks (UGNN) rey2022Untrained 7550
U: DAU-based RED (LR-Unsupervised) 22
@@ 5.1.2 Comparison Methods
Methods to be compared are summarized in Table 1. They include both of model-based and data-driven methods. Hyperparameters in the model-based methods are determined from training data using Bayesian optimization optuna_2019. The network architecture of GAT is the same as rey2022Untrained. The number of iterations $K$ of the proposed method and PnP-ADMM is set to $K=10$, considering the trade-off between denoising performance and computational cost. The internal graph denoiser $\mathcal{D}_{\text{graph}}$ for PnP-ADMM is LR. The proposed method uses two denoisers, LR or PnP-ADMM, for regularization: Their abbreviations are RED (LR) and RED (PnP), respectively. In the proposed supervised and unsupervised methods, parameters are trained using Adam DBLP:journals/corr/KingmaB14 with the learning rate $0.01$. In the proposed unsupervised method, $\sigma_{\text{N2N}}$ in (8) is randomly selected from $0,0.4\cdot\max(\mathbf{|x|})]$.
@@ 5.3 Results: Real-world Dataset
5.4 Computation Time
Table 4 shows the average computation time for the model- and RED-based methods  [^2] All experiments were performed on a workstation with an AMD Ryzen 7 9700X, NVIDIA GeForce RTX 3090, and 64 GB of memory.. The proposed methods exhibit longer runtimes compared to their baseline counterparts (LR and PnP) because they involve iterative calls to these denoisers within the RED algorithm. The increase is particularly pronounced for RED (PnP), which has a hierarchical structure: PnP iterations are nested within RED iterations, and LR is computed within PnP. We consider this increased computational cost as a trade-off for achieving higher denoising performance. As shown in Tables 2 and 3, the proposed methods significantly improve the RMSE that justify the additional runtime.
A comparison across datasets also reveals that, the execution time tends to increase for almost all methods as the number of nodes increases. This is mainly attributed to the matrix inversion of the graph Laplacian. However, this bottleneck can be mitigated by employing iterative algorithms, such as the conjugate gradient method, instead of computing the LR denoiser in a closed form.
Table 4: Average computation time of each method [ms].
Method Synthetic ($N=100$) 3-D point cloud ($N=500$)
LR 5.23 4.97
PnP 5.03 5.65
RED (LR) 6.76 19.96
RED (PnP) 27.01 284.46
RED (LR, DAU) 7.32 18.35
RED (PnP, DAU) 27.07 284.58